At the beginning of 2018, Ace Company had the following portfolio of investments in available-for-sale debt securities (all of which were acquired at par value):

Security

Cost

1/1/18 Fair Value

A $25,000 $31,000
B 38,000 36,000
Totals $63,000 $67,000

During 2018, the following transactions occurred:

May 3 Purchased C debt securities at their par value for $50,000.
July 1 Sold all of the A securities for $31,000 plus interest of $1,000.
Dec. 31 Received interest of $1,000 on the B and C securities. Additionally the following information was available:

Security

12/31/18 Fair Value

B $42,000
C 53,000

Required:

1. Prepare journal entries to record the preceding information.
2. What is the balance in the Unrealized Holding Gain/Loss account on December 31, 2018?
3. Next Level What justification does the FASB give for its treatment of unrealized holding gains and losses for available-for-sale securities?
Answer #1
W.N Calculation of Unrealized gain /loss as on 1/1/2018
Cost Fair Value-1/1/18 Unrealized Gain/(Loss)
Security A 25000 31000 6000
Security B 38000 36000 -2000
1 Prepare journal entries to record the preceding information.
Date Account Title Debit Credit
May.3, 2018 Available for sale debt securities-C $50,000
Cash $50,000
(Purchased debt securities)
July.1, 2018 Cash $32,000
Unrealized Gain - OCI $6,000
Available for sale debt securities (A) $31,000
Interest Income $1,000
Realized Gain on Sale of Debt Security A $6,000
Dec.31, 2018 Cash $1,000
Interest Income $1,000
(Interest received on B and C debt securities)
Dec.31, 2018 Available for sale debt securities (B) $6,000
Available for sale debt securities (C ) $3,000
Unrealized Gain (loss) - OCI $9,000
(Fair value adjustment as on Dec31, 2018)
WN Unrealized Gain (B) = 42000 - 36,000 = $6000
Unrealized Gain (C ) = 53,000 - 50,000 = $3,000
2 What is the balance in the Unrealized Holding Gain/Loss account on December 31, 2018?
Beginning balance of Unrealized Gain/(Loss) of Security B -2000
Add:fair value adj as on Dec 31 2018
Unrealized gain -B (42000-36000) 6000
Unrealized gain -C (53000-50000) 3000
Unrealized Gain account Balance as on Dec 31 2018 7000
3 Next Level What justification does the FASB give for its treatment of unrealized holding gains and losses for available-for-sale securities?
The justification given by FASB regarding treatment of unrealized holding gains/losses available for sale securities is that the debt
securities which are held for selling in the near future are classified as held for sale securities and thus reported at fair value.
The unrealized gains will not be realized until and unless the sale of securities is done and subsequently the cash is realized.